# RobustSpring: Benchmarking Robustness to Image Corruptions for Optical Flow, Scene Flow and Stereo

### v1

@@ Abstract
Standard benchmarks for optical flow, scene flow, and stereo vision algorithms generally focus on model accuracy rather than robustness to image corruptions like noise or rain. Hence, the resilience of models to such real-world perturbations is largely unquantified. To address this, we present RobustSpring, a comprehensive dataset and benchmark for evaluating robustness to image corruptions for optical flow, scene flow, and stereo models. RobustSpring applies 20 different image corruptions, including noise, blur, color changes, quality degradations, and weather distortions, in a time-, stereo-, and depth-consistent manner to the high-resolution Spring dataset, creating a suite of 20,000 corrupted images that reflect challenging conditions. RobustSpring enables comparisons of model robustness via a new corruption robustness metric. Integration with the Spring benchmark enables public two-axis evaluations of both accuracy and robustness. We benchmark a curated selection of initial models, observing that accurate models are not necessarily robust and that robustness varies widely by corruption type. RobustSpring is a new computer vision benchmark at spring-benchmark.org that treats robustness as a first-class citizen to foster models that combine accuracy with resilience.
Figure 1: RobustSpring is a novel image corruption benchmark for optical flow, scene flow and stereo. It evaluates 20 image corruptions including blurs, color changes, noises, quality degradations, and weather, applied to stereo video data from [43]. For comprehensive robustness evaluations on all three tasks, RobustSpring’s image corruptions are integrated in time, stereo and depth where applicable.
@@ 1 Introduction
11footnotetext: University of Stuttgart, VIS-CV. first.last@vis.uni-stuttgart.de
Optical flow, scene flow, and stereo vision algorithms estimate dense correspondences and enable real-world applications like robot naviation [40, 80, 34], video processing [44], structure-from-motion [38, 50], medical image registration [47] or surgical assistance [56, 51]. While estimation quality continuously improves on accuracy-driven benchmarks [43, 45, 12, 10, 57, 18, 55, 62, 5, 7], their robustness to real-world visual corruptions like sensor noise or compression artifacts is rarely systematically assessed. This lack of systematic assessment is problematic, as better accuracy does not necessarily translate to improved robustness and can even harm model robustness [71, 60]. Though image data in KITTI [45], Sintel [12] or Spring [43] comes with degradations like motion blurs, depth-of-field or brightness changes, they result from real-world data capture or efforts to increase data realism, but were not included to systematically study model predictions under image corruptions. Broad corruption-robustness studies as they exist for for image classification [21, 48], 3D object detection [46, 32] or monocular depth estimation [30] are rare for dense-correspondence tasks, where studies are limited to specific degradations like weather [61] or low-light [82]. This not only leaves uncertainty about the reliability of dense matching algorithms in real-world scenarios. It also prevents systematic efforts to improve their robustness.
To enable systematic studies on the image corruption robustness of optical flow, scene flow, and stereo, we propose the RobustSpring dataset. Based on Spring [43], it jointly benchmarks the robustness of all three tasks on corrupted stereo videos. While prior image corruptions affect the monocular 2D or 3D space [21, 30, 46], RobustSpring’s image corruptions are integrated in time, stereo and depth and thus tailored to dense matching tasks. A principled corruption robustness metric and public benchmark website make RobustSpring the first systematic tool to evaluate and improve dense matching robustness to image corruptions.
Contributions. Fig. 1 gives an overview of RobustSpring. In summary, we make the following contributions:
@@
    Tailored image corruptions. RobustSpring is the first image corruption dataset for optical flow, scene flow and stereo. It integrates 20 corruptions for blurs, noises, tints, artifacts, and weather in time, stereo, and depth.
@@
    Corruption robustness metric. We propose a corruption robustness metric, based on Lipschitz continuity, which subsamples the clean-corrupted prediction difference and disentangles robustness and accuracy.
@@
    Benchmark functionality. RobustSpring’s standardized evaluation enables community-driven robustness comparisons of dense matching models. Public robustness benchmarking can be integrated with Spring’s website.
@@
    Initial robustness evaluation. We benchmark eight optical flow, two scene flow and six stereo models. All models are corruption sensitive, which reveals concealed robustness deficits on dense matching models.
Intended Use. RobustSpring is not a fine-tuning dataset, but a benchmark of how dense matching models generalize to unseen image corruptions. It seeks to foster robustness research and, simultaneously, helps assess real-world applicability of models. Hence, it is essential to tie RobustSpring to an existing accuracy benchmark like Spring, as this minimizes the robustness evaluation hurdle for researchers.
@@ 2 Related Work
While the quality of optical flow, scene flow and stereo models advanced for over three decades, their robustness recently regained attention as result of brittle deep learning generalization [53, 60]. We review robustness in dense-matching, particularly image corruptions and metrics.
@@
Robustness in Dense Matching. Robustness research for optical flow, scene flow, and stereo models often focuses on adversarial attacks, which quantify prediction errors for optimized image perturbations. Most attacks are for optical flow [4, 61, 60, 63, 53, 77, 33] rather than stereo [11, 74] and scene flow [72, 37]. As remedies to adversarial vulnerability [3, 2, 1, 63, 9] may be overcome through specialized optimization [58], another line of robustness research considers unoptimized data shifts. Those come in two flavors: generalization across datasets, and robustness to image corruptions. Dense matching models typically report generalization [42, 69, 70, 36, 23, 76] to several datasets, which span synthetic [43, 12, 55, 39, 15, 17, 54, 35, 5, 7, 8] and real-world data [18, 45, 31, 57, 62], often in automotive contexts. While some datasets contain image corruptions, e.g. motion blur, depth of field, fog, noise or brightness changes [66, 12, 43, 45], they do not systematically assess corruption robustness. Yet, in the wild, robustness to image corruptions is crucial. For optical flow, systematic low light [82] and weather datasets [59, 61] exist, and [63, 78] apply 2D image corruptions [21, 7] to optical flow data. Beyond these isolated works on optical flow, no systematic image-corruption study before RobustSpring spans all three dense matching tasks and includes scene flow or stereo.
@@
Robustness to Image Corruptions. Popularized by 2D common corruptions [21], the field of image corruption robustness rapidly expanded from classification [21, 48] to depth estimation [30], 3D object detection [46, 32] and semantic segmentation [32]. Conceptually, corruptions were extended to the 3D space  [30], LiDAR [32], procedural rendering [16] and optical flow [78], but none have been tailored to the depth-, stereo-, and time-dependent setup of dense matching with optical flow, scene flow and stereo.
@@
Robustness Metrics and Benchmarks. Most robustness metrics for dense matching differ by whether they utilize ground truth [53, 4, 78] or not [60, 61, 59]. However, multiple works [60, 71, 68] evidence that robustness and accuracy are competing qualities whose quantification should not be mixed, which informs our robustness metric. RobustSpring is the first robustness benchmark for three dense-matching tasks, and joins prior benchmarks for optical flow [6], disparity estimation [5] and classification [14, 29, 67]
@@ 3 RobustSpring Dataset and Benchmark
RobustSpring is a large, novel, image corruption dataset for optical flow, scene flow, and stereo. Below, we describe how we build on Spring’s stereo video dataset and augment its frames with diverse image corruptions integrated in time, stereo, and depth, how we evaluate robustness to image corruptions, and use it to benchmark algorithm capabilities.
@@
Figure 2: RobustSpring’s image corruptions on a single image.
Brightness Fog Motion blur Pixelate Rain Speckle noise
Figure 3: RobustSpring example frames. The first row shows clean and corrupted images. The second row shows the left and right disparity maps predicted with LEA Stereo [13]. The third row shows the target disparities for forward left, backward left, forward right, and backward right directions from M-FUSE [42]. The fourth row shows optical flow estimates for forward left, backward left, forward right, and backward right from RAFT [69]. All disparities and flows are computed on the corrupted dataset, see Fig. A1 in the Supplementary for additional frames.
Spring Data. Spring [43] is a high-resolution benchmark and dataset with rendered stereo sequences. It is the ideal base for an image corruption dataset as its detailed renderings permit image alterations of varying granularity – from removing detail by blurring to adding detail via weather. Being a benchmark, Spring has a public training and closed test split, which withholds ground truth for optical flow, disparity, and extrinsic camera parameters. Because our robustness benchmark shall complement accuracy analyses, we use the 2000 Spring test frames, two per stereo camera. For image corruptions with time, stereo, and depth consistency, however, we require the extrinsic camera parameters and depths that are withheld. Thus, we estimate extrinsics using COLMAP 3.8 and depths as $Z=\frac{f_{x}\cdot B}{d}$, with focal length $f_{x}$, baseline length $B$ and stereo disparities $d$, estimated via MS-RAFT+ [27, 26]. Estimation also prevents data leakage and maintains ground truth confidentiality.
3.1 Corruption Dataset Creation
@@
Table 1: Overview of RobustSpring corruptions, whether they are implemented in a time, stereo and depth consistent manner, and their resulting visual changes w.r.t. the original images as SSIM.
RobustSpring corrupts the Spring test frames via 20 diverse image corruptions, summarized in Fig. 2 and Tab. 1. Below, we describe the image corruption types, their new consistencies, their implementation, and their severity levels.
Corruption Types. In RobustSpring, we consider the five image corruption types from [21]: color, blur, noise, quality, and weather. Color simulates different lighting conditions and camera settings, including brightness, contrast, and saturation. Blur acts like focus and motion artifacts, including defocus, Gaussian, glass, motion, and zoom blur. Noise represents sensor errors and ambiance, including Gaussian, impulse, speckle, and shot noise. Quality distortions are lossy compressions and geometric distortions, including pixelation, JPEG, and elastic transformations. Weather enacts outdoor conditions, including spatter, frost, snow, rain, and fog. All corruptions are on a single frame in Fig. 2.
Corruption Consistencies. To increase the realism of these 20 corruptions for dense matching models, we extend their definition to time, stereo, and depth: Time consistent corruptions are smooth over time on one camera, e.g. frost on a camera lens, which differs per stereo camera. Stereo consistent corruptions equally influence both stereo cameras, e.g. brightness changes affect the cameras to the same extent. Depth consistent corruptions are integrated into the 3D scene, e.g. snowflakes falling along a trajectory in the 3D space, rendered into the camera view. Tab. 1 summarizes the consistencies we added to 16 of our 20 corruptions. Note that depth-aware motion blur is not stereo-consistent because it depends on the specific camera view.
@@
Corruption Implementation. Though most corruptions are loosely based on [21], our corruption consistencies requires multiple adaptations. Furthermore, we employ specialized techniques for highly consistent effects, i.e. motion blur, elastic transform, snow, rain and fog. We adapt implementations from [21], modify glass blur, zoom blur, frost and pixelation to accommodate higher resolutions and non-square images, and adjust frost, glass blur, and spatter for consistency across video scenes. Motion blur is based on [81] and adds camera-induced motion with clean optical flow estimates. Elastic transform uses PyTorch’s transforms package to create a see-through water-like effect, changing object morphology with smooth frame transitions. For snow and rain, we expand  [61]’s two-step 3D particle rendering to multi-step particle trajectories and stereo views, change from additive-blending to order-independent alpha blending [41], and include global illumination [19]. To augment the large-scale Spring data, we improve its performance via more effective particle generation and parallel processing. Fog is based on the Koschmieder model following [73]. Full implementation details are in Sec. B.2.
@@
Corruption Severity. Prior works [21, 48, 30, 46, 32] defined corruptions with several levels of severity. Here we opt for one severity per corruption, because evaluating one scene flow model on all 20 corruptions already produces 2.1 TB of raw data – 1.2 GB after subsampling, cf. Sec. 3.2. More severity levels would overburden the evaluation resources of RobustSpring benchmark users. To balance severity across corruptions, we tune their hyperparameters until the image SSIM reaches a defined threshold. We generally use SSIM $\geq$ 0.7, and, because the SSIM is less sensitive to blurs than noises [22], SSIM $\geq$ 0.2 for noises for visually similar artifact strengths. Final SSIMs are in Tab. 1.

### v2

@@ Abstract
Standard benchmarks for optical flow, scene flow, and stereo vision algorithms generally focus on model accuracy rather than robustness to image corruptions like noise or rain. Hence, the resilience of models to such real-world perturbations is largely unquantified. To address this, we present RobustSpring, a comprehensive dataset and benchmark for evaluating robustness to image corruptions for optical flow, scene flow, and stereo models. RobustSpring applies 20 different image corruptions, including noise, blur, color changes, quality degradations, and weather distortions, in a time-, stereo-, and depth-consistent manner to the high-resolution Spring dataset, creating a suite of 20,000 corrupted images that reflect challenging conditions. RobustSpring enables comparisons of model robustness via a new corruption robustness metric. Integration with the Spring benchmark enables two-axis evaluations of both accuracy and robustness. We benchmark a curated selection of initial models, observing that robustness varies widely by corruption type, and experimentally show that evaluations on RobustSpring indicate real-world robustness. RobustSpring is a new computer vision benchmark to treat robustness as a first-class citizen, fostering models that are accurate and resilient.
Figure 1: RobustSpring is a novel image corruption benchmark for optical flow, scene flow, and stereo. It evaluates 20 image corruptions including blurs, color changes, noises, quality degradations, and weather, applied to stereo video data from Spring. For comprehensive robustness evaluations, RobustSpring’s image corruptions are integrated in time, stereo, and depth, where applicable.
@@ 1 Introduction
Optical flow, scene flow, and stereo vision algorithms estimate dense correspondences and enable real-world applications like robot navigation (McGuire et al., 2017; Zhang et al., 2025; Lamberti et al., 2024), video processing (Mehl et al., 2024), structure-from-motion (Maurer et al., 2018; Phan et al., 2020), medical image registration (Mocanu et al., 2021), or surgical assistance (Rosa et al., 2019; Philipp et al., 2022). While estimation quality continuously improves on accuracy-driven benchmarks (Mehl et al., 2023b; Menze and Geiger, 2015; Butler et al., 2012; Baker et al., 2011; Scharstein et al., 2014; Geiger et al., 2012; Richter et al., 2017; Schöps et al., 2017), their robustness to real-world visual corruptions like noise or compression artifacts is rarely systematically assessed. This lack of systematic assessment is problematic, as better accuracy does not necessarily translate to improved robustness, and can even harm model robustness (Tsipras et al., 2019; Schmalfuss et al., 2022b). Though image data in KITTI (Menze and Geiger, 2015), Sintel (Butler et al., 2012), or Spring (Mehl et al., 2023b) comes with degradations like motion blurs, depth-of-field, or brightness changes, they result from real-world data capture or efforts to increase data realism, but were not included to systematically study model predictions under image corruptions. Broad corruption-robustness studies as they exist for image classification (Hendrycks and Dietterich, 2019; Müller et al., 2023), 3D object detection (Michaelis et al., 2019; Kong et al., 2023) or monocular depth estimation (Kar et al., 2022) are rare for dense-correspondence tasks, where studies are limited to specific degradations like weather (Schmalfuss et al., 2023) or low-light (Zheng et al., 2020). This not only leaves uncertainty about the reliability of dense matching algorithms in real-world scenarios. It also prevents systematic efforts to improve their robustness.
To enable systematic studies on the image corruption robustness of optical flow, scene flow, and stereo, we propose the RobustSpring dataset. Based on Spring (Mehl et al., 2023b), it jointly benchmarks robustness of all three tasks on corrupted stereo videos. While prior image corruptions affect the monocular 2D or 3D space (Hendrycks and Dietterich, 2019; Michaelis et al., 2019; Kar et al., 2022; Schmalfuss et al., 2025), RobustSpring’s image corruptions are integrated in time, stereo, and depth and thus tailored to dense matching tasks. A principled corruption robustness metric and an accompanying benchmark framework make RobustSpring the first systematic tool to evaluate and improve dense matching robustness to image corruptions.
Contributions. Figure 1 gives an overview of RobustSpring. In summary, we make the following contributions:
@@
    Tailored image corruptions. RobustSpring is the first image corruption dataset for optical flow, scene flow, and stereo. It integrates 20 corruptions for blurs, noises, tints, artifacts, and weather in time, stereo, and depth.
@@
    Corruption robustness metric. We propose a corruption robustness metric based on Lipschitz continuity, which subsamples the clean-corrupted prediction difference and disentangles robustness and accuracy.
@@
    Benchmark functionality. RobustSpring’s standardized evaluation enables community-driven robustness comparisons of dense matching models. Public robustness benchmarking is integrated with Spring’s website (https://spring-benchmark.org/).
@@
    Initial robustness evaluation. We benchmark nine optical flow, two scene flow, and six stereo models. All models are corruption sensitive, which reveals concealed robustness deficits on dense matching models.
Intended Use. RobustSpring is not a fine-tuning dataset, but a benchmark of how dense matching models generalize to unseen image corruptions. It seeks to foster robustness research and, simultaneously, helps assess real-world applicability of models. Hence, it is essential to tie RobustSpring to an existing accuracy benchmark like Spring, as this minimizes the robustness evaluation hurdle for researchers. While RobustSpring treats corruptions as perturbations to assess robustness, their interpretation depends on the application domain. For instance, in autonomous driving, rain or snow are typically considered disturbances to be ignored for stable navigation, whereas in video editing or cinematic rendering, they may constitute meaningful scene content. To accommodate such differences, RobustSpring provides results per corruption type, allowing end-users to focus only on those corruptions that align with their intended application.
@@ 2 Related Work
While the quality of optical flow, scene flow, and stereo models advanced for over three decades, their robustness recently regained attention as result of brittle deep learning generalization (Ranjan et al., 2019; Schmalfuss et al., 2022b). We review robustness in dense-matching, particularly image corruptions and metrics.
@@
Robustness in Dense Matching. Robustness research for optical flow, scene flow, and stereo models often focuses on adversarial attacks, which quantify prediction errors for optimized image perturbations. Most attacks are for optical flow (Agnihotri et al., 2024c; Schmalfuss et al., 2023; 2022b; Schrodi et al., 2022; Ranjan et al., 2019; Yamanaka et al., 2021; Koren et al., 2022) rather than stereo (Berger et al., 2022; Wong et al., 2021) and scene flow (Wang et al., 2024a; Mahima et al., 2025). As remedies to adversarial vulnerability (Agnihotri et al., 2024b; a; 2023; Schrodi et al., 2022; Anand et al., 2020) may be overcome through specialized optimization (Scheurer et al., 2024), another line of robustness research considers non-adversarial data shifts. Those come in two flavors: generalization across datasets, i.e. the Robust Vision Challenge (http://www.robustvision.net/) or effective robustness (Bauer et al., 2025), and robustness to image corruptions. Work on dense matching models typically reports generalization (Mehl et al., 2023a; Teed and Deng, 2020; 2021; Lipson et al., 2021; Huang et al., 2022; Xu et al., 2022b) to several datasets, which span synthetic (Mehl et al., 2023b; Butler et al., 2012; Richter et al., 2017; Mayer et al., 2016; Dosovitskiy et al., 2015; Gaidon et al., 2016; Ranjan et al., 2020; Li et al., 2024) and real-world data (Geiger et al., 2012; Menze and Geiger, 2015; Kondermann et al., 2016; Scharstein et al., 2014; Schöps et al., 2017), often in automotive contexts. While some datasets contain image corruptions, e.g. motion blur, depth of field, fog, noise, or brightness changes (Sun et al., 2021; Butler et al., 2012; Mehl et al., 2023b; Menze and Geiger, 2015), they do not systematically assess corruption robustness. Yet, in the wild, robustness to image corruptions is crucial. For optical flow, systematic low light (Zheng et al., 2020) and weather datasets (Schmalfuss et al., 2022a; 2023) exist, and Schrodi et al. (2022); Yi et al. (2024) apply 2D image corruptions (Hendrycks and Dietterich, 2019) to optical flow data. Beyond these isolated works on optical flow, no systematic image-corruption study before RobustSpring spans all three dense matching tasks and includes scene flow or stereo.
@@
Robustness to Image Corruptions. Popularized by 2D common corruptions (Hendrycks and Dietterich, 2019), the field of image corruption robustness rapidly expanded from classification (Hendrycks and Dietterich, 2019; Müller et al., 2023) to depth estimation (Kar et al., 2022), 3D object detection (Michaelis et al., 2019; Kong et al., 2023), and semantic segmentation (Kong et al., 2023). Conceptually, corruptions were extended to the 3D space  (Kar et al., 2022), LiDAR (Kong et al., 2023), and procedural rendering (Drenkow and Unberath, 2024), but none have been tailored to the depth-, stereo-, and time-dependent setup of dense matching with optical flow, scene flow, and stereo.
@@
Robustness Metrics and Benchmarks. Most robustness metrics for dense matching differ by whether they utilize ground truth (Ranjan et al., 2019; Agnihotri et al., 2024c; Yi et al., 2024) or not (Schmalfuss et al., 2022b; 2023; a). However, multiple works (Schmalfuss et al., 2022b; Tsipras et al., 2019; Taori et al., 2020) show that robustness and accuracy are competing qualities that should not be quantified together. This informs our robustness metric. RobustSpring is the first dense-matching robustness benchmark, and joins prior classification robustness benchmarks (Croce et al., 2021; Jung et al., 2023; Tang et al., 2021)
@@ 3 RobustSpring Dataset and Benchmark
RobustSpring is a large, novel image corruption dataset for optical flow, scene flow, and stereo. Below, we describe how we build on Spring’s stereo video dataset and augment its frames with diverse image corruptions integrated in time, stereo, and depth, how we evaluate robustness to image corruptions, and use it to benchmark algorithm capabilities.
Spring Data. Spring (Mehl et al., 2023b) is a high-resolution benchmark with rendered stereo sequences and dense ground truth. It is the ideal base for an image corruption dataset as its detailed renderings permit image alterations of varying granularity – from removing detail by blurring to adding detail via weather. Spring provides a public training and closed test split, where test ground truth for optical flow, disparity, and extrinsic camera parameters is withheld. As RobustSpring is designed to complement accuracy analyses, we build on the 2000 Spring test frames (two per stereo camera). To apply corruptions with time, stereo, and depth consistency, we require depth and extrinsics that are not publicly available. We therefore estimate extrinsics using COLMAP 3.8 and depths via $Z=\tfrac{f_{x}\cdot B}{d}$, with focal length $f_{x}$, baseline $B$, and disparities $d$ predicted by MS-RAFT+ (Jahedi et al., 2022; 2024). This estimation avoids ground-truth leakage while maintaining benchmark integrity. Quantitative results on the accuracy of our depth and extrinsics estimation are given in App. A.5, and a detailed discussion of motion ranges in Spring is provided in App. A.7.
3.1 Corruption Dataset Creation
RobustSpring corrupts the Spring test frames via 20 diverse image corruptions, summarized in Fig. 2(a) and Fig. 2(b). Below, we describe the image corruption types, their new consistencies, their implementation, and their severity levels.
Corruption Types. In RobustSpring, we consider the five image corruption types from Hendrycks and Dietterich (2019): color, blur, noise, quality, and weather. Color simulates different lighting conditions and camera settings, including brightness, contrast, and saturation. Blur acts like focus and motion artifacts, including defocus, Gaussian, glass, motion, and zoom blur. Noise represents sensor errors and ambiance, including Gaussian, impulse, speckle, and shot noise. Quality distortions are lossy compressions and geometric distortions, including pixelation, JPEG, and elastic transformations. Weather enacts outdoor conditions, including spatter, frost, snow, rain, and fog. All corruptions applied to the same frame are shown in Fig. 2(a).
While these 20 corruptions do not cover the entire corruption space, they are chosen to represent the most common perturbations encountered in real imagery and to provide a balanced basis for robustness evaluation. Several implementations were adapted and we additionally include practically relevant corruptions such as saturation, Gaussian blur, speckle noise, rain, and spatter. The blur and noise families span the most dominant optical and sensor degradations, and the weather and quality corruptions capture major outdoor and compression-related effects. At the same time, some perturbations, such as illumination changes requiring re-rendering (e.g. colored or dynamic lighting), lie beyond what can be approximated in post-processing. These, along with further outdoor effects (e.g. bloom, glare, dusty conditions) or extended codec distortions (e.g. JPEG 2000), form natural directions for future extensions.
Corruption Consistencies. To increase the realism of these 20 corruptions for dense matching models, we extend their definition to time, stereo, and depth: Time-consistent corruptions evolve smoothly over subsequent frames for a single camera, mirroring persistent lens or sensor effect, e.g. frost follows a temporally coherent pattern on one camera but differs between left and right. Stereo-consistent corruptions equally influence both stereo cameras, such as shared brightness or contrast adjustments. Unlike simply using the same hyperparameters, stereo consistency does not imply identical pixel-wise noise realizations, only that both views undergo the same transformation strength. Depth-consistent corruptions are rendered directly in the 3D scene, ensuring that their projection into each stereo view respects geometry. This applies to weather effects, such as snow, rain, and fog, where particles follow 3D trajectories and generate view-dependent projections. Other corruptions, such as blur or noise, do not benefit from 3D rendering; therefore, they use independent realizations per frame, despite sharing global severity parameters. Fig. 2(b) summarizes the consistencies we added to 16 of our 20 corruptions. Note that motion blur is not stereo-consistent because it depends on the specific camera view.
@@ 3.1 Corruption Dataset Creation
(a) Image corruptions on a single image.
@@
(b) Overview of corruptions and their consistency in time, stereo, or depth, with resulting visual changes w.r.t. the original images as SSIM.
Figure 2: Overview of RobustSpring’s image corruptions.
@@
Corruption Implementation. Though most corruptions are loosely based on Hendrycks and Dietterich (2019), our corruption consistencies require multiple adaptations. Furthermore, we employ specialized techniques for highly consistent effects, i.e. motion blur, elastic transform, snow, rain and fog. We adapt implementations from Hendrycks and Dietterich (2019), modify glass blur, zoom blur, frost, and pixelation to accommodate higher resolutions and non-square images, and adjust frost, glass blur, and spatter for consistency across video scenes. Motion blur is based on Zheng et al. (2006) and adds camera-induced motion with clean optical flow estimates. Elastic transform uses PyTorch’s transforms package to create a see-through-water-like effect, changing object morphology with smooth frame transitions. For snow and rain, we expand the two-step 3D particle rendering of Schmalfuss et al. (2023) to multi-step particle trajectories and stereo views, change from additive-blending to order-independent alpha blending (McGuire and Bavoil, 2013), and include global illumination (Halder et al., 2019). To augment the large-scale Spring data, we improve its performance via more effective particle generation and parallel processing. Fog is based on the Koschmieder model following Wiesemann and Jiang (2016). Full details are in App. A.3.3.
@@
Corruption Severity. Prior works (Hendrycks and Dietterich, 2019; Müller et al., 2023; Kar et al., 2022; Michaelis et al., 2019; Kong et al., 2023) defined corruptions with several levels of severity. Here, we opt for one severity per corruption, because evaluating one scene flow model on all 20 corruptions already produces 2.1 TB of raw data – 1.2 GB after subsampling, cf. Sec. 3.2. More severity levels would overburden the evaluation resources of benchmark users. To balance severity across corruptions, we tune their hyperparameters until the image SSIM reaches a defined threshold. We generally use SSIM $\geq$ 0.7, and, because the SSIM is less sensitive to blurs than noises (Hore and Ziou, 2010), SSIM $\geq$ 0.2 for noises for visually similar artifact strengths. We conducted a focused perceptual study to validate the selection of corruption strengths and their corresponding SSIM values in RobustSpring. See App. A.6 for more details. Final SSIMs are in Fig. 2(b).
Brightness Fog Motion blur Pixelate Rain Speckle noise
Figure 3: RobustSpring example frames. First row shows clean and corrupted images. Second row shows the left and right disparity maps predicted with LEA Stereo (Cheng et al., 2020). Third row shows the target disparities for forward left, backward left, forward right, and backward right directions from M-FUSE (Mehl et al., 2023a). Fourth row shows optical flow estimates for forward left, backward left, forward right, and backward right from RAFT (Teed and Deng, 2020). All disparities and flows are computed on the corrupted dataset. See Fig. 6 in App. A.3.1 for additional frames.